OpenStudy (anonymous):

Which points are on the graph of f(x) = 5^x? Choose all answers that are correct. (-2, -25) (-1, 1/5) (0, 1) (2, 10)

OpenStudy (anonymous):

@jim_thompson5910 @ganeshie8

jimthompson5910 (jim_thompson5910):

how far did you get?

OpenStudy (anonymous):

Er...No where? Does f equal y?

jimthompson5910 (jim_thompson5910):

yes, y = f(x)

jimthompson5910 (jim_thompson5910):

plug in x to get y or f(x)

jimthompson5910 (jim_thompson5910):

let's say you plug in x = 2 f(x) = 5^x f(2) = 5^2 f(2) = 25 so when x = 2, y is y = 25. The point is really (2,25) and NOT (2,10)

OpenStudy (anonymous):

OK. I think I get it. I think they're suppose to be more than one answer though.

OpenStudy (anonymous):

HWA HWA HWAAAAHHH MWUSHIE WUSHIE ok never mind jk jk XD HAVE SUM FUN IN UR LIFE! XD

OpenStudy (anonymous):

@jim_thompson5910

jimthompson5910 (jim_thompson5910):

yes there is more than one answer in this case

jimthompson5910 (jim_thompson5910):

what answers do you get?

OpenStudy (anonymous):

(-2, -25) (2, 10)

jimthompson5910 (jim_thompson5910):

recall that above I said that the point should be (2,25) and NOT (2,10)

jimthompson5910 (jim_thompson5910):

since 5^2 = 5*5 = 25

OpenStudy (anonymous):

C?

jimthompson5910 (jim_thompson5910):

why C

OpenStudy (anonymous):

I'm just guessing. I see how the first one makes sense but not the others.

jimthompson5910 (jim_thompson5910):

for part c, x = 0, agreed?